# LLMind: Orchestrating AI and IoT with LLM for Complex Task Execution

### v3

@@ Abstract
The exploration of large language models (LLMs) for task planning and IoT automation has recently gained significant attention. However, existing works suffer from limitations in terms of resource accessibility, complex task planning, and efficiency. In this paper, we present LLMind, an LLM-based AI agent framework that enables effective collaboration among IoT devices for executing complex tasks. Inspired by the functional specialization theory of the brain, our framework integrates an LLM with domain-specific AI modules, enhancing its capabilities. Complex tasks, which may involve collaborations of multiple domain-specific AI modules and IoT devices, are executed through a control script generated by the LLM using a Language-Code transformation approach, which first converts language descriptions to an intermediate finite-state machine (FSM) before final precise transformation to code. Furthermore, the framework incorporates a novel experience accumulation mechanism to enhance response speed and effectiveness, allowing the framework to evolve and become progressively sophisticated through continuing user and machine interactions.
Introduction
Large Language Model (LLM) has recently attracted tremendous attention from both academia and industry. In addition to excellent language skills, LLM also exhibits a high level of reasoning and planning capability [1, 2, 3, 4, 5]. Moreover, the remarkable feature of in-context learning substantially broadens the range of potential application scenarios for LLM, sparking the design of diverse general-purpose AI agents [6, 7, 8, 9, 10, 11, 12, 13].
In particular, LLM presents new transformative opportunities for the IoT era. In this work, we propose LLMind, an AI agent framework that transforms the usage of conventional IoT devices by integrating LLM as an orchestrator for accomplishing complex tasks using them.
@@ Introduction
Compared with scripted intelligence and traditional IoT device control methods, the LLM orchestrator brings unparalleled flexibility and empowers IoT devices with advanced intelligence, enabling the coordination of diverse IoT devices to collaboratively accomplish intricate tasks. By harnessing the in-context learning capability of the LLM, the agent can achieve swift adaptation and optimal resource utilization, leading to efficient task solutions [14, 15]. Consequently, users can effortlessly express their needs without concerns about operational details.
@@
Integrating LLM into IoT device control presents a number of challenges before its promising potentials can be realized. First, a general LLM cannot handle specialized tasks like object detection or facial recognition, while there are already many AI models excelling in various specific domains. Thus, a challenge lies in combining the power of specialized but fragmented AI modules with LLM to handle various complex tasks. Furthermore, due to LLM’s slow inference speed and high computational costs, it is crucial to investigate methods that can enhance the system’s response speed and overall efficiency. Another significant challenge involves the effective control of IoT devices by the LLM-based AI agent. The LLM must accurately learn the diverse functionalities and operational characteristics of various IoT devices. Additionally, the design of efficient interaction mechanisms that enable users to naturally and intuitively express their needs to the LLM-based AI agent presents a further challenge. Our framework effectively overcomes the aforementioned challenges.
@@
First, taking inspiration from the functional specialization theory of the brain, our framework integrates domain-specific AI modules with the LLM, which serves as the task allocation governor. This synthesis enables the general-purpose LLM to effectively execute specific subtasks by invoking specialized AI modules to accomplish an overall complex task, enhancing system flexibility and practicality. Moreover, by incorporating new AI modules, the AI agent acquires new capabilities and continues to evolve to become even more capable.
@@
Second, to enhance the response speed and overall efficiency of the agent, we utilize a semantic analysis NLP model to classify and locally store LLM-generated plans (control scripts) for specific user commands. By identifying similar instructions in historical records, even with varying tones or expressions, the agent can retrieve validated solutions, thus avoiding complete replanning. This approach enables the system to accumulate experience and progressively enhance its proficiency through interactions.
@@
Third, the agent leverages the LLM for complex task planning and the generation of the corresponding control script using a novel finite-state machine approach, enabling efficient and precise invocation of AI modules and IoT devices. This methodology significantly improves accuracy and success rates.
@@
Last but not least, we connect our demo system to the widely used social media platform, WeChat, allowing users to conveniently interact with the LLM agent.
@@
We refer to the intelligence level attained by our framework as Artificial Complex Intelligence (ACI), which serves as a bridge between Artificial Narrow Intelligence (ANI) [16, 17] and Artificial General Intelligence (AGI) [18, 19, 17], as illustrated in Fig. 1. ACI integrates the capabilities of various ANI modules and undergoes continuous evolution towards AGI. Our framework goes beyond mere device control, providing a comprehensive blueprint for a future intelligent, integrated, and collaborative IoT device ecosystem empowered by LLM.
Overview of the framework design
@@
Our framework, depicted in Fig. 2, consists of five components: user interface, LLM, coordinator, AI modules, and IoT devices.
The user can communicate with the AI agent via social media platforms to provide instructions and assign tasks. Once the task is comprehended, the LLM generates a control script to orchestrate IoT devices and AI modules, facilitating the completion of the assigned task.
If an error occurs during script execution, the error message is used as feedback and temporarily stored as context for subsequent trials. The successfully executed script will be locally stored and associated with the corresponding user instructions, providing a valuable learned experience for future executions.
Upon successful execution of the script or in the case of failure after multiple attempts, the agent collects the results and generates a report for the user.
User interface of the framework
Our proposed framework is open to various interaction methods such as text, pictures, and voices. It depends on valid LLM input or the availability of methods to convert user input into a suitable format for LLM processing. Social media software is a good choice since it offers a seamless and natural interaction experience that aligns with our daily lives.
Functionalities of the LLM within the framework
The LLM in our framework performs two key functionalities.
@@
First, it engages in conversations with users, generating contextually appropriate responses. To enhance the lifelike and immersive response of the LLM, we adopt a role-playing technique based on the LLM’s in-context learning ability. In our demo system, the LLM assumes the persona of a housekeeper, seamlessly blending into our everyday scenarios.
@@
Second, the LLM does the planning and generates control scripts to address the tasks raised by the user, utilizing specialized AI modules and IoT devices as part of the available resources and methods. Notably, the LLM itself does not directly interact with AI modules and IoT devices. Instead, it generates control scripts to invoke AI modules and send control commands to IoT devices through network connections. This design enables the system to adapt to different AI modules and IoT devices without requiring extensive modifications to the LLM itself, enhancing scalability and interoperability. Python was selected as the programming language for our demo system due to its wide adoption in the field of artificial intelligence.
@@
Within this framework, we propose a Language-Code transformation scheme using a finite-state machine (FSM) as a strategic and justified step for generating scripts. The complexity of user tasks often necessitates breaking them down into manageable subtasks. The FSM acts as an intermediate step, providing a structured approach to connect and sequence these subtasks. This choice is inspired by the inherent ability of FSMs to represent complex processes through a set of states and transitions, ensuring a systematic and organized generation of control scripts.
Functionalities and composition of the coordinator
@@
The coordinator in our framework serves as the central hub, facilitating the interconnection and seamless operation of various system components. Without a central coordinator, the system components would operate in isolation, leading to inefficiencies and potential conflicts. The coordinator ensures smooth functioning of the agent system. It comprises three major parts: context repository, experience archive, and script executor.
@@
    API Descriptions: A comprehensive description of the API functions used to invoke AI modules and control IoT devices within the script, encompassing details such as inputs, outputs, specifications, expected time costs, and other relevant information. This enables the LLM to effectively leverage the available functionality.
@@
The experience archive serves as a database for previously generated scripts. Valid scripts are stored locally and linked to their corresponding user instructions. This enables the system to reuse preexisting scripts when a similar instruction is given by the user in the future, resulting in improved efficiency and faster response speed. Importantly, the agent’s ability continues to grow through the accumulation and generalization of experience.
@@
The significance of introducing AI modules
Currently, LLM cannot effectively handle many domain-specific tasks [20]. To address this limitation, we integrated domain-specific AI modules into the framework, aiming to enhance overall performance by assigning specialized tasks to dedicated modules, resulting in improved capabilities, accuracy, and performance beyond LLM’s general knowledge.
The individual AI modules can be well versed in various specialized AI techniques, such as computer vision, natural language processing, speech recognition, sentiment analysis, recommendation systems, and more [21, 22, 23, 24, 25, 26, 27, 28, 29]. Each module is trained and optimized for its specific task, enabling it to provide highly accurate and specialized results.
@@
In the context of IoT device control, AI modules play a crucial role in handling specialized tasks that may be beyond the scope of a general-purpose AI model like an LLM. For example, tasks like object detection, facial recognition, anomaly detection, or localization require specialized training data and algorithms. By incorporating dedicated AI modules for these tasks, the AI system can provide more accurate and reliable results, enhancing the overall system performance and user experience.
@@
Enabling efficient and secure control of IoT devices
IoT devices encompass a wide range of physical objects embedded with sensors, actuators, and connectivity features that enable internet connectivity and data exchange [30, 31]. Examples include smart appliances, wearables, industrial equipment, and environmental sensors.
To ensure efficient control of diverse IoT devices in our proposed framework, manufacturers must expose a set of well-documented API functions. These functions should encapsulate the necessary functionality, enabling the AI agent to initiate specific actions or retrieve information from the IoT devices.
It is important to emphasize that manufacturers should prioritize security considerations when exposing API functions for remote control [32]. By implementing robust authentication mechanisms, access controls, and encryption protocols [33, 34, 35, 36] within the API functions, they can ensure that only authorized entities have control over the devices. This approach guarantees secure and protected communications, safeguarding against unauthorized access or tampering.
Experiment configurations
@@
We deployed the coordinator and AI modules of the system on an edge server. The edge server is connected to the same local network as the IoT devices via WiFi and can interact with ChatGPT using OpenAI’s APIs [37].
@@
    Security cameras: The cameras are positioned within the room to capture images of the surrounding environment. These images are transmitted to the edge server via a local network. This experiment employs three cameras: two placed at the corners of the room, approximately 2 meters above the ground, and the third mounted on top of a TurtleBot mobile robot.
@@
    TurtleBot mobile robot [38]: The TurtleBot mobile robot can receive target position coordinates through a network connection and autonomously plan paths to navigate within the room, intelligently avoiding obstacles.
@@
Check-in and security scenario experiment
Guests arrive at a party. The user sends a message to the AI agent: “Can you help me count the number of people in the room and identify who they are?” [^1] Demo video 1: https://youtu.be/3Al5qRntEEU Table I provides the dialogue content between different roles in the experiment. Among them, the LLM acts as the housekeeper, while the coordinator serves as the assistant.
@@ 3 Assistant $\rightarrow$ Housekeeper Do you require any assistance?
In step 3, the coordinator questions the LLM about the need for assistance. Following each round of prompt and response interaction between the user and the LLM (similar to steps 1 and 2), the coordinator repeats the same question to the LLM. In the scenario where the LLM assumes the role of a remote housekeeper who relies on assistants in the user’s residence to carry out tasks. This inquiry serves to verify if the user has given specific instructions. Ambiguous instructions, as well as unrelated casual conversations, will cause the system to prompt the user to provide additional information or be filtered out. When the LLM indicates a need for help, it signifies that the user has assigned a task that could potentially be handled using AI modules and IoT devices. We employ a role-playing methodology that allows for the implementation of various roles and role-playing settings to facilitate the context-filtering process.
@@
Network management scenario experiment
Our framework is designed to handle both one-off complex tasks and scenarios that involve multiple interactions with users. Next, we consider a scenario where a family of three individuals shares the same home WiFi network. [^2] Demo video 2: https://youtu.be/aTGD8EjQ8kM The interaction process between the AI agent and the user in this scenario is similar to the one described in the check-in and security scenario.
@@
Conclusion
The integration of LLMs into IoT device control offers transformative possibilities for developing intelligent IoT ecosystems. Our framework seamlessly combines LLMs’ language skills, reasoning abilities, and in-context learning with specialized AI modules, enhancing the flexibility and practicality of IoT systems. The proposed approaches, including FSM-based control script generation, user interaction, and historical script retrieval mechanisms, contribute to enhancing user experiences and system efficiency, with the system becoming increasingly intelligent as it learns from interactions among LLMs, humans, AI modules, and IoT devices. This comprehensive blueprint for intelligent and collaborative IoT ecosystems sets the stage for future advancements, where LLMs and AI-driven solutions enhance device coordination and convenience in our daily lives harmoniously.
Methods
Workflow
The operational workflow of a demo system is illustrated in the flowchart presented in Fig. 4.
Figure 4: System operational flowchart
Users can interact with the system by sending text messages through social media software. The LLM promptly acknowledges the receipt of the message. For casual chats and greetings, the system provides text replies without further processing. However, for specific requests and commands, the system formulates plans and strives to fulfill the user’s instructions using integrated AI modules and connected IoT devices.
LLM role-playing
The LLM role-playing approach is supported by carefully crafted system-level prompts, as shown in Algorithm 1. It is important to note that in order to optimize the user experience, distinct user profiles are necessary for tailoring responses to individual users. These user-related details enable the system to provide personalized interactions, catering precisely to each user’s preferences and needs.
Algorithm 1 System-level prompts for role-playing
Role-playing prompts
1: You are a serious housekeeper managing the employer’s daily life.
5: Your reply should be simple and concise.
6: You should use the same language as your employer.
User information prompts
1: Your employer is male.
2: The name of your employer is Eason.
FSM-based Language-Code transformation using LLM
Our Language-Code transformation scheme tackles two critical issues.
Compared to the “Language-Code” transformation, our novel “Language-FSM-Code” transformation scheme maintains precision and conciseness in conveying information since an FSM can precisely and comprehensively describe a task [42, 43]. Additionally, by providing APIs that enable the LLM to reference and interact with AI modules and IoT devices, the complexity of sub-task transformation within each FSM state is manageable for current LLMs. Notably, our scheme enhances reliability without causing information loss.
Figure 5: Language-Code transformation procedure
Fig. 5 depicts the Language-Code transformation procedure, which consists of three phases: First, context expansion; next, Language-FSM transformation; and finally FSM-Code transformation. Our scheme utilizes pre-stored knowledge, including environment information, API documentation of AI modules and IoT devices, and example codes, to aid in the transformation.
For context expansion, the system expands the task description using the environment information stored in the context repository of the coordinator. This information encompasses a list of available AI modules and IoT devices, their functional descriptions, as well as IoT device locations. The environment information enables a precise and executable task description.
The Language-FSM transformation leverages the expanded task description and API information to generate the FSM representation of the task. The provision of API information helps the LLM to gain a better understanding of the capabilities of AI modules and IoT devices, thereby facilitating the development of a reasonable and efficient FSM design. Without API information, the FSM representation may include overly complex sub-tasks or redundant processes, causing inefficiencies.
Historical script retrieval mechanism
To enhance system efficiency, previously generated scripts can be reused when the user issues similar instructions in the future. However, a challenge lies in how the system evaluates the similarities between user instructions. Our system implements a semantic analysis NLP model that maps words to a vector space [44]. The semantic similarity is then determined by computing the distance between these vectors. To enhance the search speed, we vectorize the instruction associated with the script before storing it in the database.
Recent works in the context of our framework
Despite the existing research on LLM-based intelligent control systems [45, 46, 47, 6, 48, 49], this remains a new area that requires further investigation to enhance user experiences and facilitate device collaboration. This section reviews several recent works in the context of our framework.
Auto-GPT [49], an open-source application built on GPT-4, enables autonomous tasks through chained LLM thoughts. It provides GPT-based models with memory and autonomy for independent task execution and experiential learning. However, Auto-GPT has limitations [50], such as high costs due to rapid token consumption, inability to retain learning across tasks or generalize reusable functions, and limited access to resources, constraining its capabilities. Our framework is an approach that allows the LLM to access AI modules and IoT devices, enriching the available resources for the agent. Additionally, we introduce a novel historical script retrieval mechanism, which enhances system efficiency and enables the system to become more experienced and skillful over time.

### v4

@@ Abstract
Task-oriented communications are an important element in future intelligent IoT systems. Existing IoT systems, however, are limited in their capacity to handle complex tasks, particularly in their interactions with humans to accomplish these tasks. In this paper, we present LLMind, an LLM-based task-oriented AI agent framework that enables effective collaboration among IoT devices, with humans communicating high-level verbal instructions, to perform complex tasks. Inspired by the functional specialization theory of the brain, our framework integrates an LLM with domain-specific AI modules, enhancing its capabilities. Complex tasks, which may involve collaborations of multiple domain-specific AI modules and IoT devices, are executed through a control script generated by the LLM using a Language-Code transformation approach, which first converts language descriptions to an intermediate finite-state machine (FSM) before final precise transformation to code. Furthermore, the framework incorporates a novel experience accumulation mechanism to enhance response speed and effectiveness, allowing the framework to evolve and become progressively sophisticated through continuing user and machine interactions.
I Introduction
Task-oriented communications and execution framework are an important trend to exploit artificial intelligence to allow IoT systems to interact with humans to execute complex tasks [1, 2]. They enable human users to control multiple IoT devices simultaneously through various media, including text, voice, video, and virtual reality. Communications between humans and IoT devices, and among IoT devices, are indispensable in such systems. Traditional communications systems emphasize on throughput, latency, and reliability. For intelligent IoT systems, beyond these requirements, a shift is needed to also emphasize “intention” at the high level [3]. Hence, the entrance of Large Language Model (LLM).
The paper proposes and demonstrates the efficacy of a framework, LLMind, that incorporates LLM to allow humans to interact with IoT devices, and for IoT devices to communicate and collaborate with each other, to perform complex tasks. LLMind not only reaffirms LLM’s impressive linguistic proficiency and exceptional logical reasoning [4], but also showcases its ability to transform conventional IoT devices into an overall IoT system that can accomplish complex tasks through collaborations with high-level instructions from humans. In particular, in LLMind, LLM serves as an orchestrator to facilitate seamless intention-oriented communications among human and IoT entities to execute complex tasks.
@@ I Introduction
The LLM orchestrator in LLMind goes beyond rigid scripted intelligence in traditional IoT device control methods. It endows IoT devices with advanced intelligence, enabling seamless coordination among diverse devices to achieve complex tasks that were previously unattainable using conventional approaches. By harnessing the in-context learning capability of the LLM, the agent can achieve swift adaptation and optimal resource utilization, leading to efficient task solutions. Consequently, users can effortlessly express their needs without concerns about operational details.
@@
Integrating LLM into IoT device control presents a number of challenges that need to be tackled before its potentials can be realized. First, currently, a general LLM cannot effectively handle specialized tasks like object detection or facial recognition, which can already be well performed by traditional AI models in specific domains. Thus, a challenge lies in combining the power of specialized but fragmented AI modules with LLM to handle various complex tasks. Furthermore, due to LLM’s slow inference speed and high computational costs, it is crucial to investigate methods that can enhance the system’s response speed and overall efficiency. Another significant challenge involves the efficient control of IoT devices through task-oriented communications by the LLM-based AI agent. The LLM must accurately learn the diverse functionalities and operational characteristics of various IoT devices. Additionally, the design of efficient interaction mechanisms that enable users to naturally and intuitively express their needs to the LLM-based AI agent presents a further challenge. Our framework effectively overcomes the aforementioned challenges.
The contributions of our work are manifold:
@@
First, inspired by the functional specialization theory of the brain, which states that different regions of the brain are dedicated to specific cognitive functions or tasks, enabling efficient and optimized processing of information, a key feature of our framework is its use of LLM to integrate domain-specific AI modules, forming an overall intelligent system. This synthesis enables the general-purpose LLM to effectively execute specific subtasks by invoking specialized AI modules to accomplish an overall complex task, enhancing system flexibility and practicality. Moreover, by incorporating new AI modules, the AI agent acquires new capabilities and continues to evolve to become ever more capable.
@@
Second, to enhance the response speed and overall efficiency of the agent, we employ a novel method to classify and locally store control scripts (LLM-generated plans) for specific user commands. By utilizing an NLP AI model to identify similar instructions in the records, even when they have varying tones or expressions, the agent can retrieve validated solutions, thus avoiding the need for complete replanning. This approach allows the system to accumulate experience and progressively improve its proficiency through interactions.
@@
Third, the agent leverages the LLM for complex task planning and the generation of the corresponding control script using a novel finite-state machine approach, enabling efficient and precise invocation of AI modules and IoT devices. This method significantly improves accuracy and success rates.
@@
Last but not least, for ease of use, we connect our demo system to WeChat, a widely used social media platform, to allow ordinary non-tech savvy users to conveniently interact with the LLM agent.
@@
We refer to the intelligence level attained by our framework as Artificial Complex Intelligence (ACI), which serves as a bridge between Artificial Narrow Intelligence (ANI) [5] and Artificial General Intelligence (AGI) [6], as illustrated in Fig. 1. ACI integrates the capabilities of various ANI modules and undergoes continuous evolution towards AGI. In summary, our framework goes beyond mere device control and provides a comprehensive blueprint for a future task-oriented, intelligent, and collaborative IoT device ecosystem empowered by LLM.
II Framework Design
@@
Our framework, depicted in Fig. 2, consists of five components: user interface, LLM, coordinator, AI modules, and IoT devices. The operational workflow of a demo system is illustrated in the flowchart presented in Fig. 3.
Figure 3: System operational process flowchart
Users can interact with the system by sending text messages through social media software. The LLM promptly acknowledges the messages. For casual chats and greetings, the system provides text replies without further processing. However, for specific requests and commands, the system formulates plans and strives to fulfill the user’s instructions using integrated AI modules and connected IoT devices.
II-A LLM
@@
The LLM in our framework performs two key functionalities. First, it engages in conversations with users, generating contextually appropriate responses. To enhance the lifelike and immersive response of the LLM, we adopt a role-playing technique based on the LLM’s in-context learning ability. In our demo system, the LLM assumes the persona of a housekeeper, seamlessly blending into our everyday scenarios.
@@
Second, the LLM does the planning and generates control scripts for user tasks utilizing specialized AI modules and IoT devices as part of the available resources and methods. Notably, the LLM itself does not directly interact with AI modules and IoT devices. Instead, it generates control scripts to invoke AI modules and send control commands to IoT devices through network connections. This design enables the system to adapt to different AI modules and IoT devices without requiring extensive modifications to the LLM itself, enhancing scalability and interoperability. We use Python as the programming language for our demo system due to its wide adoption in the field of artificial intelligence.
II-A1 LLM role-playing
The LLM role-playing approach is supported by carefully crafted system-level prompts, as shown below.
Role-playing prompts
1: You are a serious housekeeper managing the employer’s daily life.
5: Your reply should be simple and concise.
6: You should use the same language as your employer.
User information prompts
1: Your employer is male.
2: The name of your employer is Eason.
It is important to note that in order to optimize the user experience, distinct user profiles are necessary for tailoring responses to individual users. These user-related details enable the system to provide personalized interactions, catering precisely to each user’s preferences and needs.
II-A2 FSM-based Language-Code transformation using LLM
@@
Within this framework, we propose a Language-Code transformation scheme using a finite-state machine (FSM) as a strategic and justified step for generating scripts. The complexity of user tasks often necessitates breaking them down into manageable subtasks. The FSM acts as an intermediate step, providing a structured approach to connect and sequence these subtasks. This choice is inspired by the inherent ability of FSMs to represent complex processes through a set of states and transitions, ensuring a systematic and organized generation of control scripts. Our Language-Code transformation scheme tackles two critical issues.
Compared to the “Language-Code” transformation, our novel “Language-FSM-Code” transformation scheme maintains precision and conciseness in conveying information since an FSM can precisely and comprehensively describe a task [8]. Additionally, by providing APIs that enable the LLM to reference and interact with AI modules and IoT devices, the complexity of sub-task transformation within each FSM state is manageable for current LLMs. Notably, our scheme enhances reliability without causing information loss.
Figure 4: Language-Code transformation procedure
Fig. 4 shows the Language-Code transformation procedure, involving three phases: context expansion, Language-FSM transformation, and FSM-Code transformation.
For context expansion, the system expands the task description using the environment information stored in the context repository of the coordinator. This enables a precise and executable task description.
The Language-FSM transformation uses the expanded task description and API information to generate the FSM representation of the task. Providing API information helps the LLM understand the capabilities of AI modules and IoT devices, leading to a reasonable and efficient FSM design. Without API information, the FSM representation may include complex sub-tasks or redundant processes, causing inefficiencies.
II-B Coordinator
@@
The coordinator in our framework acts as a central hub, connecting and enabling the seamless operation of system components. Without a central coordinator, the system components would operate in isolation, leading to inefficiencies and potential conflicts. The coordinator ensures smooth functioning of the agent system. It comprises three major parts: context repository, experience archive, and script executor.
@@
    API Descriptions: Comprehensive descriptions of the API functions used to invoke AI modules and control IoT devices within the script, encompassing details such as inputs, outputs, specifications, expected time costs, and other relevant information. This enables the LLM to effectively leverage the available functionality.
@@
The experience archive serves as a database for previously generated scripts. Valid scripts are stored locally and linked to their corresponding user instructions. This enables the system to reuse preexisting scripts when a similar instruction is given by the user in the future, resulting in improved efficiency and faster response speed. Importantly, the agent’s ability continues to grow through the accumulation and generalization of experience. However, a challenge lies in how the system evaluates the similarities between user instructions. Our system implements a semantic analysis NLP model that maps words to a vector space [9]. The semantic similarity is then determined by computing the distance between these vectors. To enhance the search speed, we vectorize the instruction associated with the script before storing it in the database.
@@
II-C AI Modules
Currently, LLM cannot effectively handle many domain-specific tasks. To address this limitation, we integrated domain-specific AI modules into the framework, aiming to enhance overall performance by assigning specialized tasks to dedicated modules, resulting in improved capabilities, accuracy, and performance beyond LLM’s general knowledge.
The individual AI modules can be well versed in various specialized AI techniques, such as computer vision, natural language processing, speech recognition, sentiment analysis, recommendation systems, and more. Each module is trained and optimized for its specific task, enabling it to provide highly accurate and specialized results.
@@ II-C AI Modules
II-D IoT Devices
To ensure efficient control of diverse IoT devices in our framework, manufacturers must provide well-documented API functions. These functions should encapsulate the necessary functionality, enabling the AI agent to initiate specific actions or retrieve information from the IoT devices. It is important to emphasize that manufacturers should prioritize security considerations when exposing API functions for remote control, thereby preventing unauthorized access or tampering.
II-E User Interface
Our framework is open to various interaction methods, relying on valid LLM input or methods to convert user input into a suitable format for LLM processing. Social media software is a good interface, seamlessly aligning with our daily lives and providing a natural interaction experience.
III Experiments
@@
We deployed the coordinator and AI modules of the system on an edge server. The edge server is connected to the same local network as the IoT devices via WiFi and can interact with ChatGPT using OpenAI’s APIs.
@@
    Security cameras: The cameras are positioned within the room to capture images of the surrounding environment. These images are transmitted to the edge server via a local network. Our experiment employs three cameras: two placed at the corners of the room, approximately 2 meters above the ground, and the third mounted on top of a TurtleBot mobile robot.
@@
    TurtleBot mobile robot: The TurtleBot mobile robot can receive target position coordinates through a network connection and autonomously plan paths to navigate within the room, intelligently avoiding obstacles.
@@
III-A Scenario 1: Check-in and Security
Guests arrive at a party. The user sends a message to the AI agent: “Can you help me count the number of people in the room and identify who they are?” (Demo video for scenario 1: https://youtu.be/3Al5qRntEEU) Table I provides the dialogue content between different roles in the experiment. Among them, the LLM acts as the housekeeper, while the coordinator serves as the assistant.
@@ 3 Assistant $\rightarrow$ Housekeeper Do you require any assistance?
In step 3, the coordinator questions the LLM about the need for assistance. Following each round of prompt and response interaction between the user and the LLM (similar to steps 1 and 2), the coordinator repeats the same question to the LLM. In the scenario where the LLM assumes the role of a remote housekeeper who relies on assistants in the user’s residence to carry out tasks. This inquiry serves to verify if the user has given specific instructions. Ambiguous instructions, as well as unrelated casual conversations, will cause the system to prompt the user to provide additional information or be filtered out. When the LLM indicates a need for help, it signifies that the user has assigned a task that could potentially be handled using AI modules and IoT devices.
@@
III-B Scenario 2: Network Management
Our framework is designed to handle both one-off complex tasks and scenarios that involve multiple interactions with users. Next, we consider a scenario where a family of three individuals shares the same home WiFi network (Demo video for scenario 2: https://youtu.be/aTGD8EjQ8kM). The interaction process between the AI agent and the user in this scenario is similar to the one described in the check-in and security scenario.
@@ III-B Scenario 2: Network Management
IV Recent Work in the Context of Our Framework
[1] presents a comprehensive task-oriented surveillance framework for virtual emotion informatics to tackle human-requested tasks. This framework places emphasis on the integration of mobile robots, UAVs, and virtual emotion informatics. It highlights the crucial role of task-oriented communication in future wireless networks, a concept that aligns with the focus of our paper.